# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 3.3 Extreme–value scaling
Our asymptotic analysis relies on extreme–value scales: deterministic sequences that characterize the typical magnitude of the largest winner equity and the largest loser shortfall in the limit of a large market. Recall the winner and loser index sets $\mathcal{W}_{T}$ and $\mathcal{L}_{T}$ with cardinalities $k_{n}$ and $m_{n}$, respectively. Under mild mixing we assume the aggregate winner and loser masses concentrate at linear scales:
$\left(\frac{W_{T}(\mathcal{P}_{n})}{n},\,\frac{D_{T}(\mathcal{P}_{n})}{n}\right)\xrightarrow{p}(\mu_{+},\mu_{-}).$ (24)
Here $\mu_{+},\mu_{-}>0$ summarize the average winner and loser magnitudes. We define the respective maxima as
$\big(\omega_{T},\,\Delta_{T}\big)=\left(\max_{i\in\mathcal{W}_{T}}\big(e_{T}(\mathfrak{p}_{i,T})\big)_{+},\,\max_{i\in\mathcal{L}_{T}}\big(-e_{T}(\mathfrak{p}_{i,T})\big)_{+}\right).$ (25)
A pair of deterministic, increasing sequences $\{b^{+}_{k}\}_{k\geq 1}$ and $\{b^{-}_{m}\}_{m\geq 1}$ constitute extreme–value scales if there exist constants $c_{+},c_{-}\in(0,\infty)$ such that, as $n\to\infty$,
$\left(\frac{\omega_{T}}{b^{+}_{k_{n}}},\,\frac{\Delta_{T}}{b^{-}_{m_{n}}}\right)\xrightarrow{p}(c_{+},c_{-}).$ (26)
@@
$\big(b^{+}_{k},\,b^{-}_{m}\big)=\left(F_{+}^{-1}\!\left(1-\frac{1}{k}\right),\,F_{-}^{-1}\!\left(1-\frac{1}{m}\right)\right).$ (27)
where $F_{+}$ and $F_{-}$ denote the distribution functions of winner-side positive equities and loser-side shortfalls, respectively.
@@
  • Power–law tails (Pareto/Fréchet). For distributions satisfying $\mathop{\bf Prob}\{(e_{T})_{+}>x\}\sim C_{+}x^{-\alpha_{+}}$ and $\mathop{\bf Prob}\{(-(e_{T}))_{+}>x\}\sim C_{-}x^{-\alpha_{-}}$ with $\alpha_{\pm}>0$,
@@
Extreme–value severity scales.
We refer to $\theta_{n}=\Theta(b_{n}/n)$ as the extreme–value severity scale. This choice aligns the haircut magnitude with the typical largest winner equity $b_{n}=b^{+}_{k_{n}}$ in a large market where $k_{n}=\Theta(n)$. Theorem B.3 (Appendix B) demonstrates that this scale is canonical: for any budget-balanced ADL with total severity $H=\theta_{n}D_{T}$, the PTSR scales asymptotically as $b^{+}_{k_{n}}/(\theta_{n}n)$. Consequently, parameter choices on the extreme–value scale preserve non-degenerate efficiency limits, whereas aggressive severities $\theta_{n}\gg b^{+}_{k_{n}}/n$ drive the PTSR to zero.
@@ 3.4 Fairness
Algorithmic fairness literature distinguishes between axiomatic approaches, which posit structural invariants (often yielding impossibility results [Arrow1951, KleinbergEtAl2018FairnessImpossibility]), and optimization approaches, which minimize disparate impact via convex programs [DworkEtAl2012FairnessAwareness]. ADL spans both: queue-based policies resemble discrete ranking mechanisms subject to Arrow-style impossibilities, while pro-rata policies admit continuous convex formulations. We adopt an axiomatic perspective to characterize “ideal” fairness, then relax it to optimize for robustness.
@@
  • Sybil Resistance: Splitting an account into multiple smaller positions (or merging them) should not alter the aggregate haircut. This ensures outcomes depend on total exposure, not identity or fragmentation. Queue-based rules violate this, as splitting a large winner can push parts of its equity lower in the priority stack.
  • Monotonicity (Stable Ordering): The policy should preserve the relative ordering of winners. If trader $A$ has more equity than trader $B$ pre-ADL, $A$ should also have more equity post-ADL. This prevents arbitrary re-ranking of survivors.
@@
In §6, we prove that capped pro-rata is the unique policy satisfying all three properties, whereas standard queueing fails Sybil resistance and monotonicity.
@@ 3.5 Moral hazard
Moral hazard is a generic incentive problem that arises when one party (the agent) can take actions that affect risk or payoffs, but those actions are not perfectly observable, verifiable, or contractible by the party bearing some of the downside (the principal); see, e.g. Holmstrom1982, LaffontMartimort2002. Classic examples include workers choosing effort under wage contracts, managers taking portfolio risk on behalf of investors, or insurers providing coverage to policyholders whose behavior may change once insured. Because the agent’s action is hidden, the principal cannot simultaneously provide full insurance (so that the agent’s payoff is insulated from shocks) and preserve strong incentives for the agent to take socially desirable actions: risk sharing necessarily distorts effort incentives. The central goal of the moral-hazard literature is therefore to characterize second-best contracts (typically linear or otherwise simple sharing rules) that optimally trade off incentive provision against insurance under information constraints [DuttingEtAl2023MultiAgentContracts, Carroll2015RobustLinearContracts].
In modern treatments, moral hazard also interacts with robustness and model uncertainty: the principal designs contracts that must perform well across a range of environments, counterparty types, or effort technologies, which further limits how close one can get to the idealized first-best benchmark [DuttingEtAl2023MultiAgentContracts]. In financial systems, these issues appear whenever risk is mutualized or losses are socialized across participants: for example, in insurance pools and in centrally cleared derivatives, where default funds and loss-sharing rules can weaken incentives for individual risk management. Our ADL setting can be viewed as an instance of this general template, with winners and losers in the market acting as principals who provide mutual insurance and the exchange acting as an agent choosing risk and loss-sharing policies; the negative results in Section 5 make this second-best trade-off precise in our model.
@@ 4 Severity Optimization
    Static Interpolation: $\theta_{t}$ is fixed or linearly interpolates between deficit-matching and full queue liquidation. This is simple but rigid.
@@
which depends only on $\theta_{\pi}$ and $D_{T}(\mathcal{P}_{n})$, not on how $H(\pi)$ is distributed across winners. Thus, any solvency functional that depends on $\pi$ only via $\{R_{t}(\pi)\}_{t}$ (e.g., breach frequencies, VaR/ES of $\sum_{t}R_{t}$) can be optimized over the scalar sequence $(\theta_{t})$, holding the haircut rule fixed. Conversely, fairness and revenue functionals depend on the winner-side post-ADL equities $\{(1-h_{\pi,i})e_{T}(\mathfrak{p}_{i})_{+}\}_{i}$ and are invariant to the choice of $\theta_{t}$ once the total budget $H_{t}$ is fixed. We exploit this separability in our numerics (§9) by tuning $\theta_{t}$ (severity control) and $h_{t}$ independently.
@@ 5 Negative Results
We begin with two structural limits for ADL. They formalize a moral-hazard trade-off between solvency and winner payoffs that cannot be removed by better policy design. The tension is instantaneous and zero-sum at the shock: covering aggregate deficit $D_{T}$ eventually requires shaving the profits of the best winner. We first argue that this tension can be viewed as a form of moral hazard, in the sense used in principal-agent contracts.
@@ 5.1 Impossibility of Avoiding Moral Hazard as $n\to\infty$
Consider a book of size $n$ where loser positions are i.i.d. with mean loss $\mu>0$, so aggregate deficit scales as $D_{T}\approx\mu n$. Suppose the top winner’s equity scales sub-linearly, e.g., $b_{n}\sim c\log n$ (light tails). If the exchange commits to a fixed severity $\bar{\theta}>0$ to maintain solvency, the total haircut budget is $H_{n}\approx\bar{\theta}\mu n$. As $n\to\infty$, this linear haircut cost overwhelms the logarithmic winner equity:
$e_{n}^{\text{post}}\approx(c\log n-\bar{\theta}\mu n)_{+}\to 0.$
Thus, to preserve any winner profit, severity must vanish at the rate $\theta_{n}=O(b_{n}/n)$. Constant severity policies therefore represent a form of growing moral hazard: they transfer an increasingly large share of tail risk from the exchange to the most profitable traders. We formalize this scaling in Appendix B.6.
@@
In our model, socialization is instantaneously zero-sum: every dollar of deficit $D_{T}$ covered by ADL must come from the winners’ equity. If ADL usage does not vanish as the market grows larger ($n\rightarrow\infty$), then one cannot simultaneously keep deficits small (solvency, low $D_{T}$) and preserve winner payoffs (high $\omega_{T}$). The proposition below quantifies this tradeoff in terms of extreme-value scales and can be read as a moral-hazard impossibility result in the principal–agent sense.
@@
Suppose that the extreme value scale of the winning trader’s equity, $\omega_{T}^{\pi}(\mathcal{P}_{n})$, is $b_{n}$. Then,
$\mathsf{PTSR}_{T}(\mathcal{P}_{n},\pi)=\mathop{\bf E{}}_{\pi}\left[\frac{\omega^{\pi}_{T}(\mathcal{P}_{n})}{D^{\pi}_{T}(\mathcal{P}_{n})}\right]\asymp\frac{b_{n}}{\theta_{n}\,n}$
In particular, the example above shows the order $b_{n}/n$ is tight: unless the severity vanishes at the extreme–value scale, $\theta_{n}=\Theta(b_{n}/n)$, one cannot preserve the best trader’s profits as $n\to\infty$. See §3 for the formal definition of extreme-value scales.
@@
  • Gaussian Winners (Light Tails): For $e_{i}\sim\mathcal{N}(0,1)$, the maximum scales as $b_{n}\sim\sqrt{2\log n}$. Thus, severity must vanish rapidly, $\theta_{n}\lesssim\sqrt{2\log n}/n$, to avoid autodeleveraging the top winner. This implies that the shortfall that can be covered goes to zero as $n\rightarrow\infty$.
  • Pareto Winners (Heavy Tails): For $e_{i}\sim\text{Pareto}(\alpha)$ with $\alpha>1$, the maximum scales as $b_{n}\sim n^{1/\alpha}$. Here, the constraint is looser but still vanishing as $n\rightarrow\infty$: $\theta_{n}\lesssim n^{-(1-1/\alpha)}$.
Even in the heavy-tailed case (infinite variance, $\alpha<2$), budget balance implies that a constant fraction severity $\theta$ wipes out the top winner almost surely as $n\to\infty$.
@@ 5.2 Excessive leverage guarantees large maximal trader loss
Consider a regime where loser leverage mass $\ell^{-}$ dominates winner leverage mass $\ell^{+}$. Aggregate shortfall scales with loser liability, $D_{T}\approx\ell^{-}b_{n}^{-}$, while the top winner’s equity scales as $\omega_{T}\approx\ell^{+}b_{n}^{+}$. If severity $\theta_{n}$ is fixed, the haircut budget $H_{n}=\theta_{n}D_{T}\approx\theta_{n}\ell^{-}b_{n}^{-}$ is large. Comparing budget to winner equity:
$\frac{H_{n}}{\omega_{T}}\approx\theta_{n}\frac{\ell^{-}}{\ell^{+}}\frac{b_{n}^{-}}{b_{n}^{+}}.$
If $\ell^{-}\gg\ell^{+}$, even a small severity $\theta_{n}$ creates a haircut larger than the top winner’s entire position. Thus, safe severity is throttled not just by market size but by the leverage imbalance $\ell^{+}/\ell^{-}$. See Appendix B.6 for the rigorous construction.
@@
Let $b_{n}^{+},b_{n}^{-}$ be the extreme value scales of the winning and losing traders, respectively. Furthermore, let the winner and loser side leverage masses be $\ell_{n}^{+},\ell_{n}^{-}$, respectively. Then we have the general EVT scaling:
$\mathop{\bf E{}}_{\pi}\!\left[\frac{\omega^{\pi}_{T}(\mathcal{P}_{n})}{D^{\pi}_{T}(\mathcal{P}_{n})}\right]\ \asymp\ \frac{\ell_{n}^{+}}{\ell_{n}^{-}}\cdot\frac{b_{n}^{+}}{b_{n}^{-}}\cdot\frac{1}{\theta_{n}}.$
The full proof is in Appendix B.
@@
  • High Loser Leverage ($\Lambda_{n}\ll 1$): If losers are $10\times$ more leveraged, the safe severity drops by $10\times$. In the extreme, if $\ell_{n}^{-}\sim n$ (e.g., a few massive whales) while $\ell_{n}^{+}\sim 1$ (retail winners), the required severity $\theta_{n}$ becomes vanishingly small even faster than the EV scale suggests.
@@
These results formalize the “second-best” nature of ADL: perfect insurance without incentive distortion is impossible. Proposition 5.1 implies that tail events necessarily penalize the best traders disproportionately, while Proposition 5.2 shows that unconstrained loser leverage amplifies this burden. This mirrors the tension in Central Counterparty (CCP) recovery in traditional clearing, where Variation Gains Haircutting (VGH) socializes losses among winners to avoid insolvency [CPMI_IOSCO_2014, Gregory2015]. Like VGH (and unlike assessment powers that call external capital), ADL is an ex-post haircut on unrealized profits, creating similar moral hazard channels [DuffieZhu2011, Pirrong2011]. Practitioners should therefore treat PTSR/PMR as key risk indicators, tuning dynamic severity caps to keep these ratios bounded.

### v3

@@ 3.3 Extreme–value scaling
Our asymptotic analysis relies on extreme–value scales: deterministic sequences that characterize the typical magnitude of the largest winner endowment (for haircut capacity) and the largest loser shortfall in the limit of a large market. Recall the winner and loser index sets $\mathcal{W}_{T}$ and $\mathcal{L}_{T}$ with cardinalities $k_{n}$ and $m_{n}$, respectively. Under mild mixing we assume the aggregate winner endowment and loser deficit masses concentrate at linear scales:
$\left(\frac{U_{T}(\mathcal{P}_{n})}{n},\,\frac{D_{T}(\mathcal{P}_{n})}{n}\right)\xrightarrow{p}(\mu_{+},\mu_{-}),$ (28)
where $U_{T}=\sum_{i}w_{i,T}$ is the total haircut capacity. Here $\mu_{+},\mu_{-}>0$ summarize the average winner endowment and loser deficit magnitudes. We define the respective maxima as
$\big(\upsilon_{T},\,\Delta_{T}\big)=\left(\max_{i\in\mathcal{W}_{T}}w_{i,T},\,\max_{i\in\mathcal{L}_{T}}\big(-e_{T}(\mathfrak{p}_{i,T})\big)_{+}\right),$ (29)
where $\upsilon_{T}$ is the maximum haircutable endowment (under PNL-only, this is the maximum positive PNL). A pair of deterministic, increasing sequences $\{b^{+}_{k}\}_{k\geq 1}$ and $\{b^{-}_{m}\}_{m\geq 1}$ constitute extreme–value scales if there exist constants $c_{+},c_{-}\in(0,\infty)$ such that, as $n\to\infty$,
$\left(\frac{\upsilon_{T}}{b^{+}_{k_{n}}},\,\frac{\Delta_{T}}{b^{-}_{m_{n}}}\right)\xrightarrow{p}(c_{+},c_{-}).$ (30)
@@
$\big(b^{+}_{k},\,b^{-}_{m}\big)=\left(F_{+}^{-1}\!\left(1-\frac{1}{k}\right),\,F_{-}^{-1}\!\left(1-\frac{1}{m}\right)\right),$ (31)
where $F_{+}$ and $F_{-}$ denote the distribution functions of winner-side positive endowments (under PNL-only, positive PNL) and loser-side shortfalls, respectively.
Extreme–value severity scales.
We refer to $\theta_{n}=\Theta(b_{n}/n)$ as the extreme–value severity scale. This choice aligns the haircut magnitude with the typical largest winner endowment $b_{n}=b^{+}_{k_{n}}$ (under PNL-only, the largest positive PNL) in a large market where $k_{n}=\Theta(n)$.
@@
  • Power–law tails (Pareto/Fréchet). For distributions satisfying $\mathop{\bf Prob}\{w_{T}>x\}\sim C_{+}x^{-\alpha_{+}}$ (under PNL-only, this is the distribution of positive PNL) and $\mathop{\bf Prob}\{(-e_{T})_{+}>x\}\sim C_{-}x^{-\alpha_{-}}$ with $\alpha_{\pm}>0$,
@@ 3.4 Fairness
Algorithmic fairness literature distinguishes between axiomatic approaches, which posit structural invariants (often yielding impossibility results [13, 76]), and optimization approaches, which minimize disparate impact via convex programs [51]. ADL spans both: queue-based policies resemble discrete ranking mechanisms subject to Arrow-style impossibilities, while pro-rata policies admit continuous convex formulations. We adopt an axiomatic perspective to characterize “ideal” fairness, then relax it to optimize for robustness.
@@
  • Sybil Resistance (Split Invariance): Let $x_{\pi,i}=h_{\pi,i}w_{i,T}$ denote seized endowment. For any winner $i$, any split $w_{i,T}=\sum_{a=1}^{r}z_{a}$ with the same economic state, and any budget $B$, if $\tilde{x}_{\pi,a}(B)$ denotes seizure on split children, aggregate seizure must be unchanged:
    $\sum_{a=1}^{r}\tilde{x}_{\pi,a}(B)=x_{\pi,i}(B).$
  • Monotonicity (Stable Ordering): The policy preserves winner ordering in surviving endowment:
    $w_{i,T}\geq w_{j,T}\implies(w_{i,T}-x_{\pi,i})\geq(w_{j,T}-x_{\pi,j}).$
@@
In §6, we prove that capped pro-rata is the unique policy satisfying all three properties. Appendix D makes the queue comparison explicit: Proposition D.1 shows that queue Sybil resistance depends on the ranking score (false for absolute scores, where splitting can change queue rank, and true for ratio scores that keep rank unchanged under proportional splits, such as Hyperliquid’s documented index [72]), while Proposition D.3 shows greedy queues can still violate monotonicity.
@@ 3.5 Moral hazard
Moral hazard is a generic incentive problem that arises when one party (the agent) can take actions that affect risk or payoffs, but those actions are not perfectly observable, verifiable, or contractible by the party bearing some of the downside (the principal); see, e.g. Holmström [68], Laffont and Martimort [80]. Classic examples include workers choosing effort under wage contracts, managers taking portfolio risk on behalf of investors, or insurers providing coverage to policyholders whose behavior may change once insured. Because the agent’s action is hidden, the principal cannot simultaneously provide full insurance (so that the agent’s payoff is insulated from shocks) and preserve strong incentives for the agent to take socially desirable actions: risk sharing necessarily distorts effort incentives. The central goal of the moral-hazard literature is therefore to characterize second-best contracts (typically linear or otherwise simple sharing rules) that optimally trade off incentive provision against insurance under information constraints [50, 32].
Moral hazard also couples to robustness: contracts must work across environments and counterparties, pushing outcomes away from the first-best benchmark [50]. The same friction arises whenever losses are socialized (e.g. insurance pools or centrally cleared derivatives) as risk sharing dulls the ability for individuals to manage risk efficiently. ADL fits this template: winners and losers supply mutual insurance while the exchange chooses the loss-sharing rule. The negative results in Section 5 formalize the resulting second-best trade-off.
@@ 4 Severity Optimization
    Static Interpolation: $\theta_{t}$ is fixed or linearly moves between deficit-matching $\theta_{t}^{\text{match}}=\min\{1,\ \sum_{i}e_{t,i}/D_{t}\}$ (so $H_{t}=B_{t}^{\star}$ when feasible) and a cap $\bar{\theta}$ (e.g. 1).
@@
which depends only on $\theta_{\pi}$ and $D_{T}(\mathcal{P}_{n})$, not on how $H(\pi)$ is distributed across winners. Thus, any solvency functional that depends on $\pi$ only via $\{R_{t}(\pi)\}_{t}$ (e.g. breach frequencies, VaR/ES of $\sum_{t}R_{t}$) can be optimized over the scalar sequence $(\theta_{t})$, holding the haircut rule fixed. Conversely, fairness and revenue functionals depend on the winner-side post-ADL endowments $\{(1-h_{\pi,i})w_{i,T}\}_{i}$ (and consequently post-ADL equities $e^{\prime}_{i,T}=c_{i,T}+\mathrm{PNL}_{i,T}-x_{\pi,i}$) and are invariant to the choice of $\theta_{t}$ once the total budget $H_{t}$ is fixed. We exploit this separability in our numerics (§9) by tuning $\theta_{t}$ (severity control) and $h_{t}$ independently.
@@ 5 Negative Results
We begin with two structural limits for ADL. They formalize a moral-hazard trade-off between solvency and winner payoffs that cannot be removed by better policy design. The tension is instantaneous and zero-sum at the shock: covering the aggregate deficit $D_{T}$ eventually requires shaving the profits of the best winner. We first argue that this tension can be viewed as a form of moral hazard, in the sense used in principal-agent contracts.
@@ 5.1 Impossibility of Avoiding Moral Hazard as $n\to\infty$
Consider a book of size $n$ where loser positions are i.i.d. with mean loss $\mu>0$, so aggregate deficit scales as $D_{T}\approx\mu n$. Suppose the top winner’s endowment (under PNL-only, positive PNL) scales sub-linearly, e.g. $b_{n}\sim c\log n$ (light tails). If the exchange commits to a fixed severity $\bar{\theta}>0$ to maintain solvency, the total haircut budget is $H_{n}\approx\bar{\theta}\mu n$. As $n\to\infty$, this linear haircut cost overwhelms the logarithmic winner endowment:
$w_{n}^{\text{post}}\approx(c\log n-\bar{\theta}\mu n)_{+}\to 0.$
Thus, to preserve any winner profit capacity, severity must vanish at the rate $\theta_{n}=O(b_{n}/n)$. Constant severity policies therefore represent a form of growing moral hazard: they transfer an increasingly large share of tail risk from the exchange to the most profitable traders. We formalize this scaling in Appendix C.6.
@@
In our model, socialization is instantaneously zero-sum: every dollar of deficit $D_{T}$ covered by ADL must come from the winners’ haircutable endowment. If ADL usage does not vanish as the market grows larger ($n\rightarrow\infty$), then one cannot simultaneously keep deficits small (solvency, low $D_{T}$) and preserve winner profit capacity (high $\upsilon_{T}$). The proposition below quantifies this tradeoff in terms of extreme-value scales and can be read as a moral-hazard impossibility result in the principal–agent sense.
@@
Suppose that the extreme value scale of the winning trader’s endowment, $\upsilon_{T}(\mathcal{P}_{n})$, is $b_{n}$ (under PNL-only, this is the maximum positive PNL). Then,
$\mathsf{PTSR}_{T}(\mathcal{P}_{n},\pi)=\mathop{\bf E{}}_{\pi}\left[\frac{\upsilon^{\pi}_{T}(\mathcal{P}_{n})}{D^{\pi}_{T}(\mathcal{P}_{n})}\right]\asymp\frac{b_{n}}{\theta_{n}\,n}$
In particular, the example above shows the order $b_{n}/n$ is tight: unless the severity vanishes at the extreme–value scale, $\theta_{n}=\Theta(b_{n}/n)$, one cannot preserve the best trader’s profit capacity as $n\to\infty$.
@@
  • Gaussian Winners (Light Tails): For $w_{i}\sim\mathcal{N}(0,1)$ (endowment distribution), the maximum scales as $b_{n}\sim\sqrt{2\log n}$. Thus, severity must vanish rapidly, $\theta_{n}\lesssim\sqrt{2\log n}/n$, to avoid autodeleveraging the top winner’s endowment. This implies that the shortfall that can be covered goes to zero as $n\rightarrow\infty$.
  • Pareto Winners (Heavy Tails): For $w_{i}\sim\text{Pareto}(\alpha)$ with $\alpha>1$, the maximum scales as $b_{n}\sim n^{1/\alpha}$. Here, the constraint is looser but still vanishing as $n\rightarrow\infty$: $\theta_{n}\lesssim n^{-(1-1/\alpha)}$.
Even in the heavy-tailed case (infinite variance, $\alpha<2$), budget balance implies that a constant fraction severity $\theta$ wipes out the top winner’s endowment almost surely as $n\to\infty$.
@@ 5.2 Excessive leverage guarantees large maximal trader loss
Consider a regime where loser leverage mass $\ell^{-}$ dominates winner leverage mass $\ell^{+}$. Let $b_{n}^{+}$ (resp. $b_{n}^{-}$) denote the extreme value (EV) scales of winner endowments and loser deficits, respectively. Aggregate shortfall scales with loser liability, $D_{T}\approx\ell^{-}b_{n}^{-}$, while the top winner’s endowment scales as $\upsilon_{T}\approx\ell^{+}b_{n}^{+}$. If severity $\theta_{n}$ is fixed, the haircut budget $H_{n}=\theta_{n}D_{T}\approx\theta_{n}\ell^{-}b_{n}^{-}$ is large. Comparing budget to winner endowment:
$\frac{H_{n}}{\upsilon_{T}}\approx\theta_{n}\frac{\ell^{-}}{\ell^{+}}\frac{b_{n}^{-}}{b_{n}^{+}}.$
If $\ell^{-}\gg\ell^{+}$, even a small severity $\theta_{n}$ creates a haircut larger than the top winner’s entire endowment. Thus, safe severity is throttled not just by market size but by the leverage imbalance $\ell^{+}/\ell^{-}$. See Appendix C.6 for the rigorous construction.
@@
Let $b_{n}^{+},b_{n}^{-}$ be the extreme value scales of the winning traders’ endowments and losing traders’ deficits, respectively. Furthermore, let the winner and loser side leverage masses be $\ell_{n}^{+},\ell_{n}^{-}$, respectively. Then we have the general EVT scaling:
$\mathop{\bf E{}}_{\pi}\!\left[\frac{\upsilon^{\pi}_{T}(\mathcal{P}_{n})}{D^{\pi}_{T}(\mathcal{P}_{n})}\right]\ \asymp\ \frac{\ell_{n}^{+}}{\ell_{n}^{-}}\cdot\frac{b_{n}^{+}}{b_{n}^{-}}\cdot\frac{1}{\theta_{n}},$
where $\upsilon^{\pi}_{T}$ is the maximum post-ADL endowment.
The full proof is in Appendix C.
@@
  • High Loser Leverage ($\Lambda_{n}\ll 1$): If losers are $10\times$ more leveraged, the safe severity drops by $10\times$. In the extreme, if $\ell_{n}^{-}\sim n$ (e.g. a few massive whales) while $\ell_{n}^{+}\sim 1$ (retail winners), the required severity $\theta_{n}$ becomes vanishingly small even faster than the EV scale suggests.
@@
These results formalize the “second-best” nature of ADL: perfect insurance without incentive distortion is impossible. Tail shocks necessarily haircut the top winners’ profit capacity (Proposition 5.1), and high loser leverage magnifies that hit (Proposition 5.2). This is the same trade-off seen in clearinghouses: Variation Gains Haircutting trims winner gains to keep the CCP solvent [45, 61]. ADL is the perpetual futures analogue of Variation Gain Haircutting (i.e., an ex-post haircut on unrealized profits), so the same moral-hazard tension applies [49, 96]. Under PNL-only haircuts, PTSR/PMR measure profit capacity survival (not total equity survival), playing the same role as VGH fairness/solvency gauges in CCP recovery [45, 61, 49].